OpenStudy (anonymous):

WILL FAN AND MEDALLL PLZZZ HELPPP!! Which of the following correctly uses absolute value to show the distance between –70 and 13? A) |–70 – 13| = |–83| = –83 units B) |–70 – 13| = |–83| = 83 units C) |–70 + 13| = |–57| = 57 units D) |–70 + 13| = |–57| = –57 units

OpenStudy (anonymous):

To find the distance between any two numbers, you subtract one number from another.

OpenStudy (anonymous):

In this case, your first number is -70 and your second number is 13.

OpenStudy (anonymous):

\[-70 - 13\] is the only correct subtraction expression, so you can eliminate C and D.

OpenStudy (anonymous):

Okay so u would subtract -70 - 13? @math1234

OpenStudy (anonymous):

Correct, now since you want to find the "distance" between two numbers, you take the absolute value of their difference.

OpenStudy (anonymous):

So is the answer B? @math1234

OpenStudy (anonymous):

An absolute value expression by definition is defined as follows: \[\left| \pm x \right| = x\]

OpenStudy (anonymous):

Yes, its B.

OpenStudy (anonymous):

Thanks :) Can you help me with one more question? I will give u a medal? @math1234

OpenStudy (anonymous):

Sure, please go ahead.

OpenStudy (anonymous):

Jinni opened a bank account. She deposited $87.60 into her account every month for 10 months. She used $25.60 every month to pay for music classes. After 10 months, she used 1 over 4 of the total money left in her account to go to a summer camp for musicians. What is the total amount of money Jinni spent to go to the summer camp? A) $64.00 B)$88.00 C)$155.00 D) $620.00

OpenStudy (anonymous):

@math1234

OpenStudy (anonymous):

Ok, first begin by writing the algebraic expression for this problem.

OpenStudy (anonymous):

Jinni deposited money for 10 months. In total, she must have deposited \[10 \times $87.60\]

OpenStudy (anonymous):

Okay so far its $876 right? @math1234

OpenStudy (anonymous):

However, Jinni spends $25.60 each month. So her remaining balance after 10 months is \[10 \times ($87.60 - $25.60) = 10 \times $62.00 = $620.00\]

OpenStudy (anonymous):

Yes correct. I just subtracted the monthly expenses too.

OpenStudy (anonymous):

So the answer will be D?

OpenStudy (anonymous):

Now of that $620.00 in her remaining balance, she spends a quarter of that. How much did she really spend?

OpenStudy (anonymous):

Look a more carefully at the question, she spends 1 over 4 of her remaining money.

OpenStudy (anonymous):

Okay so would u divide?

OpenStudy (anonymous):

Yes, you should divide $620.00 by 4.

OpenStudy (anonymous):

So its C? @math1234

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

Thanks for ur help :) @math1234